# ChatGPT’s One-year Anniversary: Are Open-Source Large Language Models Catching up?

### 28 Nov 2023 v1

@@ 3.6.2 Safety
Models
Based on current evaluations (Zhang et al., 2023c; Röttger et al., 2023), GPT-3.5-turbo and GPT-4 models remain at the top for safety evaluations. This is largely attributed to Reinforcement Learning with Human Feedback (RLHF) (Bai et al., 2022a). RLHF first collects a human preference dataset on responses, then trains a reward model to mimic human preferences, and finally uses RL to train the LLM to align with human preferences. In the process, LLMs learn to demonstrate desired behaviors and exclude harmful responses such as impolite or biased answers. However, the RLHF procedure requires collecting a large number of expensive human annotations, which hinders its use for open-source LLMs. To democratize endeavors on advancing the safety alignment of LLMs, Ji et al. (2023) gathers a human-preference dataset to disentangle harmlessness and helpfulness from the human-preference score, thus providing separate ranking data for the two metrics. Experiments show that disentangling human preferences enhances safety alignment. Bai et al. (2022b) seeks to increase safety with RL from AI Feedback (RLAIF), where the preference model is trained using LLM-generated self-critiques and revisions. Direct Preference Optimization (DPO) (Rafailov et al., 2023a) reduces the need to learn a reward model and learn from preferences directly with a simple cross-entropy loss, which could largely reduce costs for RLHF. Combining and improving these methodologies could lead to potential improvements in safety for open-source LLMs.
@@ 4.1 Development Trend of LLMs
Ever since Brown et al. (2020) demonstrated that a frozen GPT-3 model can achieve impressive zero- and few-shot performance on a variety of tasks, numerous efforts have been made to advance the development of LLMs. One line of research focused on scaling up model parameters, including Gopher (Rae et al., 2021), GLaM (Du et al., 2022), LaMDA (Thoppilan et al., 2022), MT-NLG (Smith et al., 2022) and PaLM (Chowdhery et al., 2022), culminating at 540B parameters. Despite exhibiting remarkable capabilities, the closed-source nature of these models limited their widespread application, thereby leading to a growing interest in developing open-source LLMs (Zhang et al., 2022; Workshop et al., 2022).
@@
Rather than scaling up model size, another line of research explored better strategies or objectives for pre-training smaller models, such as Chinchilla (Hoffmann et al., 2022b) and UL2 (Tay et al., 2022). Beyond pre-training, considerable attention has been devoted to studying instruction tuning of LMs, e.g., FLAN (Wei et al., 2021b), T0 (Sanh et al., 2021) and Flan-T5 (Chung et al., 2022).
@@
The emergence of OpenAI’s ChatGPT a year ago greatly changed the research focus of the NLP community (Qin et al., 2023a). To catch up with OpenAI, Google and Anthropic introduced Bard and Claude, respectively. While they show comparable performance to ChatGPT on many tasks, there is still a performance gap between them and the latest OpenAI model GPT-4 (OpenAI, 2023b). As the success of these models is primarily attributed to reinforcement learning from human feedback (RLHF) (Schulman et al., 2017b; Ouyang et al., 2022a), researchers have explored various ways to improve RLHF (Yuan et al., 2023; Rafailov et al., 2023b; Lee et al., 2023b).
@@
To promote research of open-source LLMs, Meta released Llama series models (Touvron et al., 2023a, b). Since then, open-source models based on Llama have started to emerge explosively. One representative research direction is to fine-tune Llama with instruction data, including Alpaca (Taori et al., 2023), Vicuna (Chiang et al., 2023), Lima (Zhou et al., 2023b) and WizardLM (Xu et al., 2023a). The ongoing research has also explored improving the agent (Xu et al., 2023d; Zeng et al., 2023; Patil et al., 2023; Qin et al., 2023b), logical reasoning (Roziere et al., 2023; Luo et al., 2023a, c) and long-context modeling (Tworkowski et al., 2023; Xiong et al., 2023; Xu et al., 2023b) capabilities of Llama-based open-source LLMs. Besides, rather than developing LLMs based on Llama, many efforts have been devoted to training powerful LLMs from scratch, e.g., MPT (Team, 2023), Falcon (Almazrouei et al., 2023), XGen (Nijkamp et al., 2023), Phi (Gunasekar et al., 2023; Li et al., 2023e), Baichuan (Yang et al., 2023a), Mistral (Jiang et al., 2023a), Grok (xAI, 2023) and Yi (01ai, 2023). We believe that developing more powerful and efficient open-source LLMs to democratize the capabilities of closed-source LLMs should be a quite promising future direction.
@@ 4.2 Summary of Results
There are also several domains where open-source LLMs are able to beat GPT-3.5-turbo. For LLM-based agents, open-source LLMs are able to surpass GPT-3.5-turbo with more extensive and task-specific pre-training and fine-tuning. For example, Lemur-70B-chat (Xu et al., 2023d) performs better in exploring the environment and following feedback on coding tasks. AgentTuning (Zeng et al., 2023) improves on unseen agent tasks. ToolLLama (Qin et al., 2023b) can better grasp tool usage. Gorilla (Patil et al., 2023) outperforms GPT-4 on writing API calls. For logical reasoning, WizardCoder (Luo et al., 2023c) and WizardMath (Luo et al., 2023a) improve reasoning abilities with enhanced instruction tuning. Lemur (Xu et al., 2023d) and Phi (Gunasekar et al., 2023; Li et al., 2023e) achieve stronger abilities by pre-training on data with higher quality. For modelling long contexts, Llama-2-long (Xiong et al., 2023) can improve on selected benchmarks by pre-training with longer tokens and a larger context window. Xu et al. (2023b) improves over 7 long-context tasks by combining context window extension with positional interpolation and retrieval augmentation. For application-specific capabilities, InstructRetro (Wang et al., 2023a) improves on open-ended QA by pre-training with retrieval and instruction tuning. With task-specific fine-tuning, MentaLlama-chat-13B (Yang et al., 2023c) outperforms GPT-3.5-turbo in mental health analysis datasets. Radiology-Llama2 (Liu et al., 2023) can improve performance on radiology reports. Stru-Bench (Tang et al., 2023b), a fine-tuned 7B model, can improve structured response generation compared to GPT-3.5-turbo, which is a core ability to support agentic tasks. Shepherd (Wang et al., 2023c), with only 7B parameters, can achieve comparable or better performance compared to GPT-3.5-turbo in generating model feedbacks and critiques. For trustworthy AI, hallucinations can be reduced by fine-tuning with data of higher quality (Lee et al., 2023a), context-aware decoding techniques (Dhuliawala et al., 2023), external knowledge augmentation such as Li et al. (2023c); Yu et al. (2023b); Peng et al. (2023); Feng et al. (2023), or multi-agent dialogue (Cohen et al., 2023; Du et al., 2023).
@@ 4.3 Recipe of Best Open-source LLMs
While the majority of LLMs utilize the decoder-only transformer architecture, different techniques in the model are employed to optimize efficiency. Llama-2 implements Ghost attention for improved multi-turn dialogue control (Touvron et al., 2023b). Mistral (Jiang et al., 2023b) employs sliding window attention to handle extended context lengths.
@@
The process of supervised fine-tuning (SFT) with instruction tuning data is vital. For high quality outcomes, tens of thousands of SFT annotations are sufficient, as evidenced by the 27,540 annotations used for Llama-2 (Touvron et al., 2023b). The diversity and quality of these data are essential (Xu et al., 2023a). In the RLHF stage, proximal policy optimization (PPO) (Schulman et al., 2017a) is often the preferred algorithm to better align the model’s behavior with human preferences and instruction adherence, playing a key role in enhancing LLM safety. An alternative to PPO is direct preference optimization (DPO) (Rafailov et al., 2023a). Zephyr-7B (Tunstall et al., 2023), for instance, employs distilled DPO and has shown results comparable to 70B-LLMs on various general benchmarks, even surpassing GPT-3.5-turbo on AlpacaEval.
@@ 4.4 Loopholes and potential problems
The issue of data contamination has become increasingly pronounced with the release of foundation models that obscure the source of their pre-training corpus. This lack of transparency can result in biased perceptions regarding the genuine generalization capabilities of Large Language Models (LLMs). Ignoring the cases where benchmark data is manually integrated into the training set with annotations from human experts or larger models, the root of the data contamination problem lies in the fact that the collecting source of benchmark data is already encompassed in the pre-training corpus. While these models are not intentionally pre-trained using supervised data, they can still acquire exact knowledge. Consequently, it is crucial to address the challenge of detecting the pre-training corpus of LLMs (Shi et al., 2023), exploring the overlap between existing benchmarks and widely-used pre-training corpus, and assessing overfitting to benchmarks (Wei et al., 2023). These efforts are essential for enhancing the faithfulness and reliability of LLMs. Looking ahead, future directions could involve establishing standardized practices for disclosing pre-training corpus details and developing methods to mitigate data contamination throughout the model development lifecycle.
@@
The application of Reinforcement Learning from Human Feedback (RLHF) for alignment using general preference data has obtained increasing attention within the community. However, only a limited number of open-source LLMs have been augmented with RLHF for alignment, primarily due to the scarcity of high-quality, publicly available preference datasets and pre-trained reward models. Some initiatives (Bai et al., 2022a; Wu et al., 2023; Cui et al., 2023) have sought to contribute to the open-source community. Yet, we are still facing the challenges lacking diverse, high-quality and scalable preference data in complex reasoning, programming, and safety scenarios.

### 15 Jan 2024 v4

@@ 3.6.2 Safety
Performance of LLMs
Based on current evaluation (Zhang et al., 2023c; Röttger et al., 2023), GPT-3.5-turbo and GPT-4 remain at the top for safety evaluations. This can be largely attributed to Reinforcement Learning with Human Feedback (RLHF) (Bai et al., 2022a), which first collects a human preference dataset on responses, then trains a reward model to mimic human preferences, and finally uses RL to train the LLM to align with human preferences. In the process, LLMs learn to demonstrate desired behaviors and exclude harmful responses such as impolite or biased answers. However, the RLHF procedure requires collecting a large number of expensive human annotations, which hinders its use for open-source LLMs. To democratize endeavors on advancing the safety alignment of LLMs, Ji et al. (2023) gather a human-preference dataset to disentangle harmlessness and helpfulness from the human-preference score, thus providing separate ranking data for the two metrics. Experiments show that disentangling human preferences enhances safety alignment. Bai et al. (2022b) seek to increase safety with RL from AI Feedback (RLAIF), where the preference model is trained using LLM-generated self-critiques and revisions. Direct Preference Optimization (DPO) (Rafailov et al., 2023a) reduces the need to learn a reward model and learn from preferences directly with a simple cross-entropy loss. Combining and improving these methodologies could lead to potential improvements in safety for open-source LLMs.
@@ 4.1 Development Trend of LLMs
Ever since Brown et al. (2020) demonstrated that a frozen GPT-3 model can achieve impressive zero- and few-shot performance on a variety of tasks, numerous efforts have been made to advance the development of LLMs. One line of research focused on scaling up model parameters, including Gopher (Rae et al., 2021), GLaM (Du et al., 2022), LaMDA (Thoppilan et al., 2022), MT-NLG (Smith et al., 2022) and PaLM (Chowdhery et al., 2022), culminating at 540B parameters. Despite exhibiting remarkable capabilities, the closed-source nature of these models limited their widespread application, thereby leading to a growing interest in developing open-source LLMs (Zhang et al., 2022; Workshop et al., 2022; Zhao et al., 2023b).
@@
Rather than scaling up model size, another line of research explored better strategies or objectives for pre-training smaller models, such as Chinchilla (Hoffmann et al., 2022) and UL2 (Tay et al., 2022). Beyond pre-training, considerable attention has been devoted to studying instruction tuning of LMs, e.g., FLAN (Wei et al., 2021), T0 (Sanh et al., 2021) and Flan-T5 (Chung et al., 2022).
@@
The emergence of OpenAI’s ChatGPT a year ago greatly changed the research focus of the NLP community (Qin et al., 2023a). To catch up with OpenAI, Google and Anthropic introduced Bard and Claude, respectively. While they show comparable performance to ChatGPT on many tasks, there is still a performance gap between them and the latest OpenAI model GPT-4 (OpenAI, 2023b). As the success of these models is primarily attributed to reinforcement learning from human feedback (RLHF) (Schulman et al., 2017; Ouyang et al., 2022a), researchers have explored various ways to improve RLHF (Yuan et al., 2023; Rafailov et al., 2023b; Lee et al., 2023b).
To promote research of open-source LLMs, Meta released Llama series models (Touvron et al., 2023a, b). Since then, open-source models based on Llama have started to emerge explosively, and some of them have achieved better performance than Llama-2-70B with much fewer parameters, e.g., Yi-34B  [^8] https://github.com/01-ai/Yi, Skywork (Wei et al., 2023b), Aquila-2  [^9] https://github.com/FlagAI-Open/Aquila2, QWen (Bai et al., 2023a), InternLM (Team, 2023a), Yuan-2.0  [^10] https://github.com/IEIT-Yuan/Yuan-2.0, XVERSE  [^11] https://github.com/xverse-ai/XVERSE-65B, Mistral (Jiang et al., 2023), and Mixtral-MoE (Jiang et al., 2024). Although some of them do not release the technical details for training, which should not be viewed as fully open-sourced, the publicly available weights can still serve as the stable foundation for future research. Besides, one another representative research direction is to fine-tune Llama with instruction data. For example, Alpaca (Taori et al., 2023) is fine-tuned from Llama on 52K instruction-following demonstrations generated by self-instruct (Wang et al., 2022b) from OpenAI’s text-davinci-003. Vicuna (Chiang et al., 2023) is an open-source chatbot fine-tuned from Llama on 70K user-shared conversations collected from ShareGPT. Lima (Zhou et al., 2023a) demonstrates that a small number of carefully annotated instruction samples can be enough to reveal the potential capabilities of LLMs fully. By leveraging Evol-Instruct to use LLMs to generate large amounts of instruction data with varying complexity, WizardLM (Xu et al., 2023a) achieves comparable performance to OpenAI’s ChatGPT.
@@
The ongoing research has also explored improving the agent (Xu et al., 2023d; Zeng et al., 2023; Patil et al., 2023; Qin et al., 2023b), logical reasoning (Roziere et al., 2023; Luo et al., 2023a, c) and long-context modeling (Tworkowski et al., 2023; Xiong et al., 2023; Xu et al., 2023b) capabilities of Llama-based open-source LLMs. Besides, rather than developing LLMs based on Llama, many efforts have been devoted to training powerful LLMs from scratch, including generic LLMs (e.g., MPT (Team, 2023b), Falcon (Almazrouei et al., 2023), XGen (Nijkamp et al., 2023), Phi (Gunasekar et al., 2023; Li et al., 2023f), Baichuan (Yang et al., 2023a), Mistral (Jiang et al., 2023), Grok (xAI, 2023) and Yi (01ai, 2023)) and domain-specific LLMs (e.g., Bloomberggpt (Wu et al., 2023a), FinGPT (Yang et al., 2023b) and InstructRetro (Wang et al., 2023b)). We believe that developing more powerful and efficient open-source LLMs to democratize the capabilities of closed-source LLMs should be a quite promising future direction.
@@ 4.2 Summary of Results
There are also several domains where open-source LLMs are able to beat GPT-3.5-turbo. For LLM-based agents, open-source LLMs are able to surpass GPT-3.5-turbo with more extensive and task-specific pre-training and fine-tuning. For example, Lemur-70B-chat (Xu et al., 2023d) performs better in exploring the environment and following feedback on coding tasks. AgentTuning (Zeng et al., 2023) improves on unseen agent tasks. ToolLLama (Qin et al., 2023b) can better grasp tool usage. Gorilla (Patil et al., 2023) outperforms GPT-4 on writing API calls. For logical reasoning, WizardCoder (Luo et al., 2023c) and WizardMath (Luo et al., 2023a) improve reasoning abilities with enhanced instruction tuning. Lemur (Xu et al., 2023d) and Phi (Gunasekar et al., 2023; Li et al., 2023f) achieve stronger abilities by pre-training on data with higher quality. For modelling long contexts, Llama-2-long (Xiong et al., 2023) can improve on selected benchmarks by pre-training with longer tokens and a larger context window. Xu et al. (2023b) improves over 7 long-context tasks by combining context window extension with positional interpolation and retrieval augmentation. For application-specific capabilities, InstructRetro (Wang et al., 2023a) improves on open-ended QA by pre-training with retrieval and instruction tuning. With task-specific fine-tuning, MentaLlama-chat-13B (Yang et al., 2023d) outperforms GPT-3.5-turbo in mental health analysis datasets. Radiology-Llama2 (Liu et al., 2023b) can improve performance on radiology reports. Stru-Bench (Tang et al., 2023b), a fine-tuned 7B model, can improve structured response generation compared to GPT-3.5-turbo, which is a core ability to support agentic tasks. Shepherd (Wang et al., 2023e), with only 7B parameters, can achieve comparable or better performance compared to GPT-3.5-turbo in generating model feedbacks and critiques. For trustworthy AI, hallucinations can be reduced by fine-tuning with data of higher quality (Lee et al., 2023a), context-aware decoding techniques (Dhuliawala et al., 2023), external knowledge augmentation such as Li et al. (2023d); Yu et al. (2023b); Peng et al. (2023); Feng et al. (2023), or multi-agent dialogue (Cohen et al., 2023; Du et al., 2023).
@@ 4.3 Recipe of Best Open-source LLMs
While the majority of LLMs utilize the decoder-only transformer architecture, different techniques in the model are employed to optimize efficiency. Llama-2 implements Ghost attention for improved multi-turn dialogue control (Touvron et al., 2023b). Mistral (Jiang et al., 2023) employs sliding window attention to handle extended context lengths.
@@
The process of supervised fine-tuning (SFT) with instruction tuning data is vital. For high quality outcomes, tens of thousands of SFT annotations are sufficient, as evidenced by the 27,540 annotations used for Llama-2 (Touvron et al., 2023b). The diversity and quality of these data are essential (Xu et al., 2023a). In the RLHF stage, proximal policy optimization (PPO) (Schulman et al., 2017) is often the preferred algorithm to better align the model’s behavior with human preferences and instruction adherence, playing a key role in enhancing LLM safety. An alternative to PPO is direct preference optimization (DPO) (Rafailov et al., 2023a). Zephyr-7B (Tunstall et al., 2023), for instance, employs distilled DPO and has shown results comparable to 70B-LLMs on various general benchmarks, even surpassing GPT-3.5-turbo on AlpacaEval.
Serving (Inference)
Inference in LLMs can be slow. New methods have been created to expedite this process. One such technique, speculative decoding (Leviathan et al., 2023; Chen et al., 2023b), accelerates sampling from autoregressive models. By leveraging approximate model generation and processing multiple tokens simultaneously, this approach can speed up the inference by 2 to 3 times without altering the results. Additionally, vLLM (Kwon et al., 2023) is a high-speed, user-friendly library specifically designed for LLM fast inference. It stands out for its state-of-the-art serving capacity and effective memory utilization.
@@ 4.4 Loopholes and potential problems
The issue of data contamination has become increasingly pronounced with the release of foundation models that obscure the source of their pre-training corpus. This lack of transparency can result in biased perceptions regarding the genuine generalization capabilities of Large Language Models (LLMs). Ignoring the cases where benchmark data is manually integrated into the training set with annotations from human experts or larger models, the root of the data contamination problem lies in the fact that the collecting source of benchmark data is already encompassed in the pre-training corpus. While these models are not intentionally pre-trained using supervised data, they can still acquire exact knowledge. Consequently, it is crucial to address the challenge of detecting the pre-training corpus of LLMs (Shi et al., 2023), exploring the overlap between existing benchmarks and widely-used pre-training corpus, and assessing overfitting to benchmarks (Wei et al., 2023a). These efforts are essential for enhancing the faithfulness and reliability of LLMs. Looking ahead, future directions could involve establishing standardized practices for disclosing pre-training corpus details and developing methods to mitigate data contamination throughout the model development lifecycle.
@@
The application of Reinforcement Learning from Human Feedback (RLHF) for alignment using general preference data has obtained increasing attention within the community. However, only a limited number of open-source LLMs have been augmented with RLHF or alike (DPO) for alignment, primarily due to the scarcity of high-quality, publicly available preference datasets and pre-trained reward models. Some initiatives (Bai et al., 2022a; Wu et al., 2023b; Cui et al., 2023) have sought to contribute to the open-source community. Yet, we are still facing the challenges lacking diverse, high-quality and scalable preference data in complex reasoning, programming, and safety scenarios.